# CARE: a Benchmark Suite for the Classification and Retrieval of Enzymes

### v1

@@ Abstract
Enzymes are important proteins that catalyze chemical reactions. In recent years, machine learning methods have emerged to predict enzyme function from sequence; however, there are no standardized benchmarks to evaluate these methods. We introduce CARE, a benchmark and dataset suite for the Classification And Retrieval of Enzymes (CARE). CARE centers on two tasks: (1) classification of a protein sequence by its enzyme commission (EC) number and (2) retrieval of an EC number given a chemical reaction. For each task, we design train-test splits to evaluate different kinds of out-of-distribution generalization that are relevant to real use cases. For the classification task, we provide baselines for state-of-the-art methods. Because the retrieval task has not been previously formalized, we propose a method called Contrastive Reaction-EnzymE Pretraining (CREEP) as one of the first baselines for this task. CARE is available at https://github.com/jsunn-y/CARE/.
@@ 1 Introduction
Proteins, which are sequences of amino acid building blocks, are not only integral components of all living organisms, but also important for a myriad of commercial applications spanning from the health domain to the bio-economy. Enzymes are a subclass of proteins that can catalyze chemical reactions, and they have many applications in areas such as bioremediation, plastic degradation, gene editing, and drug synthesis [1, 2, 3].
@@
Identifying the specific chemical reactions that an enzyme is capable of performing (i.e. the enzyme’s function) is a key first step for many applications [4]. For instance, during standard metagenomic analyses, enzyme genes are annotated for their functions, which enables determination of an organism’s critical metabolic pathways and specialization [5, 6, 7]. When applied to chemical synthesis, enzyme annotations are needed to identify catalysts that can replace existing step(s) in drug synthesis procedures–known as retrobiosynthesis [8]. Likewise, enzyme engineering for new-to-nature function [9] involves discovering an enzyme starting point with a desired function before improving activity for that function using protein engineering techniques [10, 11]. Historically, similarity search algorithms, most notably BLAST, have been the most common methods used to assign function to protein sequences [12, 13]. These methods work by finding similar sequences in annotated reference databases, as similar sequences are likely to share function. Still, there is a need for more abundant, high-quality annotations of enzymes and automated workflows to identify enzymes with desired functions.
@@
In recent years, there has been increasing interest in using machine learning (ML) for a broad range of applications related to the functional prediction and design of enzymes [11, 14, 15]. In particular, ML models have emerged to classify protein sequences based on their function, which are reviewed here [16] with some more recent examples listed here [17, 7, 18, 19, 20, 21, 22]. Despite these advancements, there is no standard benchmark or dataset for evaluating computational models for enzyme function prediction. A challenge associated with classification is that a given enzyme is often able to perform multiple reactions [23], and many reactions are not annotated. Moreover, complex tasks such as extrapolation to unannotated reactions have yet to be evaluated.
@@
In this work, we present a benchmark suite for the Classification And Retrieval of Enzymes (CARE). Our contributions can be summarized as: (1) formalizing model evaluation into two tasks which are directly relevant to scientists and engineers: classification of an enzyme by function (Task 1), and retrieval of enzyme sequences based on a reaction (Task 2); (2) curating high-quality and easy-to-use datasets; (3) providing train-test splits that mimic challenging extrapolations in real-world use cases; and (4) benchmarking state-of-the-art models for Task 1, and providing the first baselines for Task 2. Because Task 2 has not been previously explored, we introduce a model called Contrastive Reaction-EnzymE Pretraining (CREEP) to serve as an initial baseline. CREEP can perform contrastive learning across three different modalities (protein, reaction, and textual description), and the learned representations are then used for retrieval. Overall, we anticipate that CARE will be a useful and easy-to-use resource for ML researchers to benchmark their enzyme function prediction models.
@@ 2 Related Work
Datasets. Various databases have emerged to help researchers store, share, and identify functionally annotated enzymes. Protein sequence databases such as UniProt [24] and Pfam [25] are catalogues of annotated protein sequences. While most databases reference protein sequences, increasingly, these sequences can be linked to protein structures, either experimentally validated, as in the Protein Data Bank (PDB), or via structural prediction tools [26] in databases such as the AlphaFold Database [27]. BRENDA [28] is a curated database specific for reaction and enzyme sequence information. Rhea [29] consolidates information from BRENDA, and other sources, such as pathway databases including KEGG [30]. There is ongoing work to compile and clean/standardize reactions from multiple databases, namely ECReact [31] and EnzymeMap [32]. Related to these databases, RetroBioCat is a database and retrobiosynthesis planning tool [33, 34], and Selenzyme is tool to retrieve enzymes to perform a target reaction [35].
@@
Protein Benchmarks. Our work takes inspiration from existing protein fitness prediction benchmarks, where fitness is a quantification of some function. TAPE [36] and FLIP [37] evaluate representations from protein language models for the prediction of a broad range of general and specific protein properties (stability, secondary structure, binding etc.) ProteinGym considers sequence variant effect prediction by using likelihoods from these language models [38]. While benchmarks for protein fitness prediction tasks are well defined, there are no standardized benchmarks for protein function prediction, which is more qualitative/categorical rather than quantitative. Here, we focus on proteins that perform chemical reactions (i.e., enzymes).
@@
Classification of Enzyme Function. ML models have emerged to predict the outcomes of enzymatic reactions [39, 40, 41] and for classification of enzyme function. Enzyme function is usually expressed using enzyme commission (EC) numbers, which is a hierarchical scheme for classifying enzyme function into classes (families) and consists of four levels of descriptions (Figure 1A). Some classification models are general protein function prediction models, which encompass all proteins, not just enzymes, such as ProtCNN/ENN [20] and ProteInfer [21]. Many models utilize representations from protein language models [42, 43, 44], and others incorporate protein structure as information, such as DeepFRI [19] and BioCLIP [45]. Methods related to supervised contrastive learning [46] have been particularly useful here, such as CLEAN and HiFi-NN [18, 7], likely by reducing imbalances in the number of sequences representing each EC number. Recently, combining ML with BLAST has enabled retrieval-augmented function prediction in ProtEx [22].
@@
Large Language Models. Recently, there has been an explosion in pretrained models in the biological and chemical domains, particularly large language models (LLMs) [47]. For example, ChatGPT is capable of answering questions related to general scientific knowledge. These language models can be further finetuned for applications such as answering questions about protein sequences (Pika, [48]) and for reaction synthesis planning (ChemCrow, [49]), among others. LLMs present important benchmarks for enzyme functional classification and retrieval, given the widespread adoption of LLMs as science facilitators and their ease of use (in particular webserver-based approaches), compared to domain-specific methods. In this study, we include ChatGPT as a baseline.
@@ 3 Overview of CARE
To streamline benchmarking, we curate a dataset of enzymes, reactions, and their associated EC numbers for CARE. At a high level, we build two datasets, one that links protein sequence to EC and one that links reaction to EC (Figure 1A). The former is processed from Swiss-Prot, the expert-validated portion of UniProt [24] and filtered to protein sequences between length 100 and 650 with annotated EC number(s). The latter is formed as a combination of EnzymeMap [32] and ECReact [31], where ECReact is only used to supplement EC numbers that are missing in EnzymeMap. Our workflow for generating the datasets used in this work is explained in detail in Appendix A.1 and shown visually in Appendix Figure A.1.
@@ 4 Task 1: Enzyme Classification
Task 1, classification of an enzyme sequence, tests the ability of a model to extrapolate to unseen protein sequences. Task 1 is a fairly well studied task [16], but model evaluation has not been previously standardized. Task 1 applies to use cases where a scientist is given an unannotated enzyme sequence and seeks to understand the enzymatic function associated with that sequence (Figure 1C). For example, an environmental biologist using metagenomic data may seek to understand which microbes in a community are utilizing energy from alkane degradation, or a retrobiosynthesis planner may look for unannotated proteins to improve a synthesis workflow. With the emergence of conditional generative models for protein sequences, it is also important to have high-throughput computational methods that can predict the function of generated sequences [58, 59, 60, 61]. For this task, a query protein is passed through a trained model to predict an EC number. The EC number is then associated with likely reactions that the protein will be able to perform.
@@
<30% Identity Held-out sequences with approximately < 30% identity to the training set 168,682 175
30-50% Identity Held-out sequences with approximately 30-50% identity to the training set 168,682 196
50-70% Identity Held-out sequences with approximately 50-70% identity to the training set 168,682 204
70-90% Identity Held-out sequences with approximately 70-90% identity to the training set 168,682 206
Previously Misclassified (Price) Previously misclassified enzyme sequences from Price et al. 168,682 146
Promiscuous Held-out sequences with multiple EC numbers 168,682 179
@@
  • <30%, 30-50%, 50-70%, and 70-90% identity splits: these four test splits contain sequences with sequence identities falling in the respective range, to sequences in the training set. Sequence identity is related to the normalized Levenshtein distance between two sequences. Natural sequences with high (>50%) sequence identity at the protein sequence level are likely to share function. It is expected that the lower the sequence identity, the more difficult it is to assign functional annotation.
  • previously misclassified (Price et al.) [5] split: challenging to assign because some have low sequence identity to other proteins, and many may lie near activity cliffs, where function can change sharply in sequence space.
@@
For Task 1, the training split is constructed by holding out all of the pooled sequences in the test splits. The classification output vocabulary (EC numbers) is closed rather than open. We verify that sequences in the test sets generally fall within the expected sequence identities, relative to the training set (Figure 2A). Notably, the Price test set has a wide distribution of sequence identities, while the promiscuous test set has high sequence similarity to the training set. The sequences in the test splits are distributed generally evenly across all different EC numbers (Figure A.4 in Appendix). More details can be found in Appendix A.2.
Task 1 benchmarking results. Benchmarking results for Task 1 are summarized in Table 2. We start with two methods as baselines, classification using a random order of EC numbers (Random), and npysearch (a python implementation of BLAST at the protein sequence level, herein referred to as BLASTp), which is a workhorse bioinformatics tool which performs local-alignment to determine the most similar sequence(s) given a target query and a database [12].
Table 2: Performance of various methods on Task 1. Performance is measured as k=1 classification accuracy (%)–except for the promiscuous test set which uses k=5 and shows the average accuracy across all true EC numbers. Similarly, for BLASTp, the top k= 1 or 5 hits were used to map to EC number(s). The 50-70% and 70-90% identity splits are omitted here because performance is very high across the board. *denotes that the model was not retrained, so there may be data leakage causing performance to be inflated. Bolded accuracy is the best model without potential data leakage.
@@
<30% Identity Random 0.0 0.6 4.0 22.3
BLASTp 52.6 56.6 57.1 65.1
ChatGPT* 0.0 0.0 2.9 16.0
ProteInfer* 61.1 65.7 68.6 72.6
CLEAN 57.7 68.6 73.7 80.6
30-50% identity Random 0.0 0.0 2.0 23.5
BLASTp 74.0 79.4 81.9 86.3
ChatGPT* 0.0 0.0 3.9 14.2
ProteInfer* 76.5 80.4 83.8 87.3
CLEAN 82.1 91.8 95.4 99.0
Previously Random 0.0 3.4 6.8 13.7
Misclassified BLASTp 41.8 71.2 79.5 81.5
(Price) ChatGPT* 0.0 3.4 6.8 13.7
ProteInfer* 13.7 47.3 56.2 67.1
CLEAN 34.2 78.1 89.0 92.5
Promiscuous Random 0.3 7.0 18.3 63.5
BLASTp 94.9 95.2 95.2 97.5
ChatGPT* 1.1 8.3 21.4 47.6
ProteInfer* 76.3 82.2 85.1 89.2
CLEAN 97.0 98.0 98.3 99.4
While there exist many ML tools to perform EC number classification, CLEAN [18], HiFi-NN [7], and ProtEx [22] seem to report the current state-of-the-art. HiFi-NN and ProtEx are not included in this study, as they are not publicly available yet. CLEAN generally performs better than BLASTp across the board, but there are certain cases where the performance is similar. We additionally benchmark ProteInfer [21], a neural network classification tool, but its performance is likely inflated as reported here, as we were unable to retrain it for this study to ensure no data leakage from the test set. ChatGPT was also tested, but it appeared to often hallucinate as it was forced to provide an answer, with results similar to the random baseline. Models that are soon to be released, such as Pika, may bridge the gap between standard LLMs and enzyme classification [48].
Overall, these results suggest that there is still room to improve classification of protein sequences with low sequence identity (<30%) and that lie near multiple activity peaks (Price). The promiscuous split is not easy for all methods, even though the test sequences have high sequence identity to the train set. Interestingly, BLAST performs close to the state-of-the-art, even when compared to more complex ML models. Future models could take advantage of this finding to augment training, as ProtEx does [22]. Additional details on the implementation of each method can be found in Appendix A.4.

### v2

@@ Abstract
Enzymes are important proteins that catalyze chemical reactions. In recent years, machine learning methods have emerged to predict enzyme function from sequence; however, there are no standardized benchmarks to evaluate these methods. We introduce CARE, a benchmark and dataset suite for the Classification And Retrieval of Enzymes (CARE). CARE centers on two tasks: (1) classification of a protein sequence by its enzyme commission (EC) number and (2) retrieval of an EC number given a chemical reaction. For each task, we design train-test splits to evaluate different kinds of out-of-distribution generalization that are relevant to real use cases. For the classification task, we provide baselines for state-of-the-art methods. Because the retrieval task has not been previously formalized, we propose a method called Contrastive Reaction-EnzymE Pretraining (CREEP) as one of the first baselines for this task and compare it to the recent method, CLIPZyme. CARE is available at https://github.com/jsunn-y/CARE/.
@@ 1 Introduction
Proteins, which are sequences of amino acid building blocks, are not only integral components of all living organisms, but also important for a myriad of commercial applications spanning from the health domain to the bio-economy. Enzymes are a subclass of proteins that can catalyze chemical reactions, and they have many applications in areas such as bioremediation, plastic degradation, gene editing, and drug synthesis [1, 2, 3, 4].
@@
Identifying the specific chemical reactions that an enzyme is capable of performing (i.e. the enzyme’s function) is a key first step for many applications [5]. While hundreds of millions of proteins have been sequenced, less than 1% are annotated with function [6]. For instance, during standard metagenomic analyses, enzyme genes are annotated for their functions, which enables determination of an organism’s critical metabolic pathways and specialization [7, 8, 9]. When applied to chemical synthesis, enzyme annotations are needed to identify catalysts that can replace existing step(s) in drug synthesis procedures–known as retrobiosynthesis [10]. Likewise, enzyme engineering for new-to-nature function [11] involves discovering an enzyme starting point with a desired function before improving activity for that function using protein engineering techniques [12, 13]. Historically, similarity search algorithms, most notably BLAST, have been the most common methods used to assign function to protein sequences [14, 15]. These methods work by finding similar sequences in annotated reference databases, as similar sequences are likely to share function. However, up to one third of genes from bacterial genomes cannot be assigned function using existing methods [7]. As such, there is a need for more abundant, high-quality annotations of enzymes and automated workflows to identify enzymes with desired functions.
@@
In recent years, there has been increasing interest in using machine learning (ML) for a broad range of applications related to the functional prediction and design of enzymes [13, 16, 17]. In particular, ML models have emerged to classify protein sequences based on their function, which are reviewed here [18] with a few examples listed here [19, 9, 20, 21, 22, 23, 24]. Despite these advances, there is no standard benchmark or dataset for evaluating computational models for enzyme function prediction [25]. A challenge associated with classification is that a given enzyme is often able to perform multiple reactions [26], and many reactions are not annotated. Moreover, complex tasks such as extrapolation to unannotated reactions [27] have yet to be evaluated.
@@
In this work, we present a benchmark suite for the Classification And Retrieval of Enzymes (CARE). Our contributions can be summarized as: (1) formalizing model evaluation into two tasks that encompass applications relevant to scientists and engineers: classification of an enzyme by function (Task 1), and retrieval of enzyme sequences based on a reaction (Task 2); (2) curating high-quality and easy-to-use datasets; (3) providing train-test splits that mimic challenging extrapolations in real-world use cases; and (4) benchmarking state-of-the-art models for Task 1, and providing a new method that serves as a baseline for Task 2. Because Task 2 has been minimally explored, we introduce a model called Contrastive Reaction-EnzymE Pretraining (CREEP) to serve as a baseline for text, reaction, and sequence integration and compare it to an existing approach (CLIPZyme) that uses sequence and reaction, in addition to other retrieval approaches. CREEP can perform contrastive learning across three different modalities (protein, reaction, and textual description), and the learned representations are then used for retrieval. Overall, we anticipate that CARE will be a useful and easy-to-use resource for ML researchers to benchmark their enzyme function prediction models.
@@ 2 Related Work
Datasets. Various databases have emerged to help researchers store, share, and identify functionally annotated enzymes. Protein sequence databases such as UniProt [6] and Pfam [28] are catalogs of annotated protein sequences. While most databases reference protein sequences, increasingly, these sequences can be linked to protein structures, either experimentally validated, as in the Protein Data Bank (PDB), or via structural prediction tools [29] in databases such as the AlphaFold Database [30]. BRENDA [31] is a curated database specific for reaction and enzyme sequence information. Rhea [32] consolidates information from BRENDA, and other sources, such as pathway databases including KEGG [33]. There is ongoing work to compile and clean/standardize reactions from multiple databases, namely ECReact [34] and EnzymeMap [35]. Related to these databases, there are retrobiosynthesis planning tools [36, 37, 38], and Selenzyme is tool to retrieve enzymes to perform a target reaction [39].
@@
Protein Benchmarks. Our work takes inspiration from existing protein fitness prediction benchmarks, where fitness is a quantification of some function. TAPE [40] and FLIP [41] evaluate representations from protein language models for the prediction of a broad range of general and specific protein properties (stability, secondary structure, binding etc.) ProteinGym considers sequence variant effect prediction by using likelihoods from these language models [42]. While benchmarks for protein fitness prediction tasks are well defined, there are no standardized benchmarks for protein function prediction. Fitness is a numerical quantification of protein function (stability, enzyme activity level, etc.), but function is more qualitative/categorical, e.g. a description of an enzymatic reaction associated with a protein. Here, we focus on proteins that perform chemical reactions, enzymes.
@@
Classification of Enzyme Function. ML models have emerged to predict the outcomes of enzymatic reactions [43, 44, 45] and for classification of enzyme function. Enzyme function is usually expressed using enzyme commission (EC) numbers, which is a hierarchical scheme for classifying enzyme function into classes (families) and consists of four levels of descriptions (Figure 1A). Some classification models are general protein function prediction models, which encompass all proteins, not just enzymes, such as ProtCNN/ENN [22] and ProteInfer [23]. Many models utilize representations from protein language models [46, 47, 48], and others incorporate protein structure as information, such as DeepFRI [21] and BioCLIP [49]. Methods related to supervised contrastive learning [50] have been particularly useful here, such as CLEAN, HiFi-NN, and Enzhier [20, 9, 51], likely by reducing imbalances in the number of sequences representing each EC number. Recently, enhanced approaches have enabled function prediction in ProtEx[24], PhiGnet [52], and others [53, 54, 55]. Other retrieval tools enable more sensitive detection of homologs such as DHR [56], ProtTrek [57], and using structure [58, 59, 60]. Finally, there are related models that predict substrates for enzymes [61, 62] and that aim to learn connections between chemical space and protein space [63].
@@
Large Language Models. Recently, there has been an explosion in pretrained models in the biological and chemical domains, particularly large language models (LLMs) [64, 65]. For example, ChatGPT is capable of answering questions related to general scientific knowledge. These language models can be further finetuned for applications such as answering questions about protein sequences (including Pika [66], InstructProtein [67], InstructBioMol [68], ProteinGPT [69], and ProteinChat [70]) and for reaction synthesis planning (ChemCrow [71]), among others. LLMs present important benchmarks for enzyme functional classification and retrieval, given the widespread adoption of LLMs as science facilitators and their ease of use (in particular webserver-based approaches), compared to domain-specific methods [72].
@@ 3 Overview of CARE
To streamline benchmarking, we curate a dataset of enzymes, reactions, and their associated EC numbers for CARE. At a high level, we build two datasets, one that links protein sequence to EC and one that links reaction to EC (Figure 1A). The former is processed from Swiss-Prot, the validated portion of UniProt [the_UniProt_consortium_UniProt_2023] and filtered to protein sequences between length 100 and 1024 with annotated EC number(s). The latter is formed as a combination of EnzymeMap [35] and ECReact [34], where ECReact is only used to supplement EC numbers that are missing in EnzymeMap. Our workflow for generating the datasets used in this work is explained in detail in Appendix A.1 and shown visually in Appendix Figure A.1.
@@ 4 Task 1: Enzyme Classification
Task 1, classification of an enzyme sequence, tests the ability of a model to extrapolate to unseen protein sequences. Task 1 is a fairly well studied task [18], but model evaluation has not been previously standardized as there are many factors to consider, such as the distribution of sequences and functions in the test sets. Task 1 applies to use cases where a scientist is given an unannotated enzyme sequence and seeks to understand the enzymatic function associated with that sequence (Figure 1C), for example for metagenomic analysis or finding new enzymes for retrobiosynthesis. With the emergence of conditional generative models for protein sequences, it is also important to have high-throughput computational methods that can predict the function of generated sequences [80, 81, 82, 83, 84]. For this task, a query protein is passed through a trained model to predict an EC number. The EC number is then associated with likely reactions that the protein will be able to perform.
@@
<30% Identity Held-out sequences with approximately < 30% identity to the training set 184,529 432
30-50% Identity Held-out sequences with approximately 30-50% identity to the training set 184,529 560
Previously Misclassified (Price) Previously misclassified enzyme sequences from Price et al. 184,529 148
Promiscuous Held-out sequences with multiple EC numbers 184,529 209
@@
  • <30% and 30-50% identity splits: these two test splits contain sequences with sequence identities falling in the respective range, to sequences in the training set by using clustering (Section A.2). Sequence identity is related to the normalized Levenshtein distance between two sequences. Natural sequences with high (>40%) sequence identity at the protein sequence level are likely to share function [86]. It is expected that the lower the sequence identity, the more difficult it is to assign functional annotation.
  • previously misclassified (Price et al.) [7] split: challenging to assign because some have low sequence identity to other proteins, and many may lie near "activity cliffs" (Figure A.3), a region where function can change sharply in sequence space.
@@
For Task 1, the training split is constructed by holding out all of the pooled sequences in the test splits. The classification output vocabulary (EC numbers) is closed rather than open, so EC numbers in the test sets are present in the train sets. We verify that sequences in the test sets generally fall within the expected sequence identities, relative to the training set (Figure 2A). Notably, the Price test set has a wide distribution of sequence identities, while the promiscuous test set has high sequence similarity to the training set. The sequences in the test splits are distributed generally evenly across all different EC numbers (Figure A.4 in Appendix). More details can be found in Appendix A.2.
Task 1 benchmarking results. Benchmarking results for Task 1 are summarized in Table 2. We start with two methods as baselines, classification using a random order of EC numbers (Random), and Diamond BLAST at the protein sequence level, herein referred to as BLASTp, which is a workhorse bioinformatics tool that performs local-alignment to determine the most similar sequence(s) given a target query and a database [14].
Table 2: Performance of various methods on Task 1. Performance is measured as k=1 classification accuracy (%). However, for the promiscuous split, we use k=(number of true ECs) for Random and CLEAN and report the average accuracy across all true EC numbers. For this split, we did not evaluate ChatGPT and Pika, as we prompted for only a single EC in the response. For BLAST we use all EC numbers associated to the top single sequence hit. More details are provided in Section A.4. CLEAN is a state-of-the-art method at the time of publication. Other methods such as ProteInfer, HiFi-NN, Enzhier, ProtEx, PhiGnet, etc. could also be benchmarked here.
@@
<30% Identity Random 0.0 1.2 3.2 19.4
BLASTp 51.4 60.0 62.5 65.7
ChatGPT 0.0 0.0 1.6 28.9
Pika 20.6 37.7 46.1 61.6
CLEAN 55.1 68.8 74.8 84.5
30-50% identity Random 0.0 0.7 3.6 22.5
BLASTp 81.1 87.9 90.7 92.3
ChatGPT 0.0 1.4 3.0 34.8
Pika 37.7 50.2 60.0 73.8
CLEAN 80.2 88.0 91.6 95.5
Previously Random 0.0 0.7 4.7 22.3
Misclassified BLASTp 35.1 70.9 78.4 78.4
(Price) ChatGPT 0.0 9.5 17.6 37.2
Pika 4.1 50.7 64.9 82.4
CLEAN 31.8 74.3 81.8 85.8
Promiscuous Random 0.5 4.1 9.0 41.1
BLASTp 93.7 94.8 95.2 95.9
CLEAN 69.4 77.9 81.5 87.0
While there exist many ML tools to directly perform EC number classification, CLEAN [20] and several others [51, 9, 24, 52, 55] are a few that seem to report the current state-of-the-art with comparable performance. Some of the latter methods were not yet publicly available at the time of this study, and other recent retrieval methods would also be promising approaches to test, but we opted to focus on benchmarking on CLEAN here. CLEAN generally performs better than BLASTp across the board, but there are certain cases where the performance is similar. ChatGPT was also tested, but it appeared to often hallucinate as it was forced to provide an answer, with results similar to the random baseline. Pika seems to bridge the gap between standard LLMs and enzyme classification, but its performance is not as good as standard tools like BLASTp [66].
Overall, these results suggest that there is still room to improve classification of protein sequences with low sequence identity (<30%) and that lie near multiple activity peaks (Price). The promiscuous split is not easy for all methods, even though the test sequences have high sequence identity to the train set – CLEAN sometimes misses EC numbers. Interestingly, BLAST performs close to the state-of-the-art, even when compared to more complex ML models. Future models could take advantage of this finding to augment training, as ProtEx does [24]. Additional details on the implementation of each method can be found in Appendix A.4.